Unable to use the IR remote control today

I have set up a new Raspberry Pi 3 with the latest OSMC (2017.07-1), and I had it working with the remote control. I left the room for several hours, and when I came back the screen was black and the machine unresponsive, so I rebooted it. After reboot, everything seemed fine, except the remote no longer worked with Kodi.

If I stopped the mediacenter service, and tested the remote with the “mode2 -d /dev/lirc0” command, I was getting the usual series of space and pulse signals OK. And in My OSMC, the correct conf file was selected for the remote.

I compared the remote against my old Pi 2, and it was continuing to work there, so the remote itself was still issuing the same sequences.

I’m not sure how to proceed to investigate this further.

While I am not a lirc expert I assume enabling debug logging in Kodi and providing the logs should help

https://osmc.tv/wiki/general/how-to-submit-a-useful-support-request/

Thanks for your reply. Unfortunately after posting this I noticed more severe errors than just the remote control, so I reinstalled OSMC, and the problem is gone. :slight_smile: